# Investigating Large Language Models for Code Vulnerability Detection: An Experimental Study

### v1

@@ 6.4 Analysis on Analysis on Datasets with Varying Sample Lengths
In general, sample length has some influence on fine-tuning LLMs. We discover that as sample length increases, the F1-score of the fine-tuned LLM decreases, though this trend is less pronounced than the effect of positive sample ratio discussed in Section 6.3. It can be concluded that positive sample ratio has a much greater impact on fine-tuning the LLM than sample length.
Figure 5: Metrics on Varying Code Sequence Length.
6.5 Sensitivity Study
@@
Figure 7: Sensitivity Study on LoRA Scaling Factor
7 Discussion & Conclusion
@@
For future works, as our analysis indicates, class imbalance is one of key factors for this task. The label noise issue also matters as discussed in Section 5.2. We aim to investigate the data quality assessment and robust training techniques tailored for the CVD task referring to [44, 45, 1, 47, 46], and evaluate the performance of more LLMs with larger parameter space to study scaling laws. Furthermore, we aim to enhance the detection performance of LLMs with the assistance of informative clues [17, 48], pre-training technique [6] and more continuously updating high-quality dataset [53] or more balanced data generation [66]. If there are some available high-quality and well-curated data, we reckon other effective post-training techniques like direct preference optimization (DPO) [67] are expected to further enhance the LLMs’ detection precision to identify vulnerable codes, which calls for more joint efforts in future.
[…]
This work is supported by the National Key Research and Development Program of China (2021YFB2900102) and the National Natural Science Foundation of China (No. 62072436 and No. 62472410).

### v2

@@ VI-D Analysis on Analysis on Datasets with Varying Sample Lengths
In general, sample length has some influence on fine-tuning LLMs. We discover that as sample length increases, the F1-score of the fine-tuned LLM decreases, though this trend is less pronounced than the effect of positive sample ratio discussed in Section VI-C. It can be concluded that positive sample ratio has a much greater impact on fine-tuning the LLM than sample length.
VI-E Sensitivity Study
@@
VII Discussion & Conclusion
@@
For future works, as our analysis indicates, class imbalance is one of key factors for this task. The label noise issue also matters as discussed in Section V-B. We aim to investigate the data quality assessment and robust training techniques tailored for the CVD task referring to [44, 45, 1, 48, 46, 47], and evaluate the performance of more LLMs with larger parameter space and different architectures (e.g. Deepseek series [77] and Mistral series [78]) to study scaling laws. Furthermore, we aim to enhance the detection performance of LLMs with the assistance of informative clues [17, 49], pre-training technique [6] and more continuously updating high-quality dataset [54] or more balanced data generation and training techniques [67, 74, 75, 76]. If there are some available high-quality and well-curated data, we reckon other effective post-training techniques like direct preference optimization (DPO) [68] are expected to further enhance the LLMs’ detection precision to identify vulnerable codes, which calls for more joint efforts in future.
VIII Acknowledgemnets
This work is supported by the National Key Research and Development Program of China (2021YFB2900102) and the National Natural Science Foundation of China (No. 62472410).